Joy cuts lawns in the summer season. Let q be the number of acres mowed per day and L be the number of hours worked per day. Joy never works more than 8 hours per day and during that time his short-run production function is q=0.2L. What is the average product of joy? How many arces can be mowed in one day?
1
Expert's answer
2019-10-18T09:37:43-0400

The average product of joy is: AP = q/L = 0.2L/L = 0.2 acres.

0.2*8 = 1.6 acres can be mowed in one day.